Let X be a random variable with the following probability distribution:





x 1 2 3 4




P(X) 0.4 0.3 0.2 0.1




Find the expected value and variance.

1
Expert's answer
2022-03-04T05:59:13-0500
"E(X)=1(0.4)+2(0.3)+3(0.2)+4(0.1)=2"

"E(X^2)=1^2(0.4)+2^2(0.3)+3^2(0.2)+4^2(0.1)=5"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2=5-(2)^2=1"